OpenStudy (nicole143):

I have a question that is asking me to pick a graph that would best help me find the solution to an equation. Can you please explain this to me? I'll put put the equation if it would help.. @mathmale @BTaylor

OpenStudy (nicole143):

*put up

OpenStudy (nicole143):

\[\frac{ 2 }{ x } - \frac{ 5 }{ x^2 } = -4\]

OpenStudy (btaylor):

I would separate this into two equations: \[y=\frac{2}{x}\] and \[y=\frac{5}{x^2}-4\] Graph those, then the intersection point is the solution.

OpenStudy (mathmale):

Nicole, great to hear from y ou. But I'm in the midst of helping someone else with a lengthy problem. I have plenty of faith in BTaylor and am confident that he can help you further.

OpenStudy (nicole143):

I don't think I'm to find the solution of the equation it's but if thats what I have to find the graph then I'll be right on that. Okay MM, talk to you later!

OpenStudy (nicole143):

* itself

OpenStudy (btaylor):

If you have a graph with both of those equations shown, then that is the one to pick.

OpenStudy (nicole143):

Okay, one moment while I work it out.

OpenStudy (nicole143):

I did the first one with the simple -2 through 2 as x to find the points of y but non of them match..

OpenStudy (btaylor):

It's not going to be an integer. What form is the answer supposed to be in?

OpenStudy (nicole143):

A graph

OpenStudy (btaylor):

That you draw or choose?

OpenStudy (nicole143):

I did the fill in for x, to find y, to then graph those points

OpenStudy (nicole143):

Choose

OpenStudy (btaylor):

hmmm...so you want one with a horizontal asymptote at 0 and one with it at -4; both vertical asymptotes will be 0

OpenStudy (nicole143):

Um, a bit confused but if what you are saying is what I think you are then I don't have one like that.. All the graphs pass through those points.. I may not be getting this right though

OpenStudy (btaylor):

Do you have a pic of the graphs?

OpenStudy (nicole143):

Yes, one second..

OpenStudy (mathmale):

Nicole: I'd suggest we stop for a moment and review what the goal of this problem is. BT was correct in suggesting that you draw both graphs on ONE set of axes; you then look for the point of intersection. Why?

OpenStudy (btaylor):

OK, this is taking a different approach than what I suggested. Now, I'd look plug in y=0 and solve for x. That will give you your x-intercepts, from which you can choose the graph. Sorry for my previously misleading instructions.

OpenStudy (nicole143):

Well i started to do the y=2/x but when I saw that non of them matched I stopped.. Do you want me to look at the other one?

OpenStudy (btaylor):

no, with the original equation.

OpenStudy (nicole143):

Oh, okay. Um, can you give me an explain of the y=0 and solving for x?

OpenStudy (btaylor):

again, slight change in plans. You are given \[\frac{2}{x}-\frac{5}{x^2}=-4\]You want to find the x-value that will give -4.

OpenStudy (btaylor):

So you solve for x. From that, look at that x-value(s) on the graphs you have. Where they intersect y=-4 are your solutions. If all the x-values match, that graph is your solution.

OpenStudy (nicole143):

Okay, can you still give me an explain? Maybe with a different problem so I can still do this one on my own?

OpenStudy (btaylor):

Sorry, I gotta go now. @mathmale can you step in please?

OpenStudy (nicole143):

Okay, thank you! @BTaylor

OpenStudy (mathmale):

BT: Yes. Thanks very much!

OpenStudy (nicole143):

So we meet again (:

OpenStudy (mathmale):

Nicole, I'm going to help you. Before we resume work, please tell me, in your own words, what you think the end goal of this math problem is.

OpenStudy (nicole143):

I think it is to find a graph that could possible tell you information about the equation with out you having to do any work but I may be wrong..

OpenStudy (mathmale):

So you're given an equation and are asked to solve it graphically, not analytically. BT's approach was a good one. He suggested that the solution would be the point at which the 2 different graphs he mentioned intersected.

OpenStudy (nicole143):

But non of my graphs intersect..

OpenStudy (mathmale):

Shoot. If you're not in a hurry, would you agree to resume this conversation tomorrow? I've been on OpenStudy for hours and am no longer thinking very clearly. Right: none intersect, BUT (I assume) all but one of them would not be helpful to you in finding a solution.

OpenStudy (mathmale):

What's your schedule like?

OpenStudy (nicole143):

I'm trying to get as much of my work done as I can before I fall asleep but I can skip this one and come back to it tomorrow if need be.

OpenStudy (mathmale):

That'd be better for me. When I can no longer concentrate, this math is no longer fun!

OpenStudy (nicole143):

Ha okay. Well hopefully I'll caught you tomorrow! Goodnight MM.

OpenStudy (mathmale):

Good morning, Nicole! What a difference eight hours of rest makes! This homework problem poses a challenge in that we're to chose which of four graphs best represents the given function\[\frac{ 2 }{ x }-\frac{ 5 }{ x^2 }=-4\] None of the four graphs represent all three terms of this equation. However, if we stop to think about it, we can tell that the graphs do not show 2/x, but only 5/x^2 or a combination of that with -4. How do we know that? Because both halves of 1/(x^2) points either upward or downward, whereas the left half of 1/x points downward and the right half points upward. See my drawing, below: